OpenStudy (anonymous):

Integral question I have to integrate this function (20x - 5x^2) / (x^2 + 9) between 4 and 0. The final answer has to be 20ln(5/3) + 15tan-1(4/3) -20. However, my calculator, even after somehow simplifying it a bit gives this: ( 40ln(5/3) -30tan-1(3/4) +15pi -40) / 2. As you can see there is something wrong with arctan integration, can anybody help or know how to simplify this with some identity?

OpenStudy (anonymous):

I'm asking for this because this is one of my practise exams, and in the exam i will have to use this calculator, and if i get a similar answer again, i will not have a clue of how to simplify it?

hero (hero):

Simplify \(\dfrac{20 - 5x^2}{x^2 + 9}\) first

OpenStudy (anonymous):

I know, partial fractions et all, but in the exam I don't think I would have time for that... ;/

hero (hero):

\(-\dfrac{5x^2 - 20}{x^2 + 9} = -\dfrac{5x^2 + 45 - 65}{x^2 + 9} = -\dfrac{5x^2 + 45}{x^2 + 9} + \dfrac{65}{x^2 + 9} = -5 + \dfrac{65}{x^2 + 9}\)

hero (hero):

You don't have to know partial fractions to do this

hero (hero):

You know that \(5(x^2 + 9) = 5x^2 + 45\)

hero (hero):

Use that to help you simplify it

hero (hero):

Now you can easily integrate this: \[\int -5 dx + \int \frac{65}{x^2 + 9} dx\]

OpenStudy (anonymous):

actually, i came across this at the other forums tan−1(3/4) +tan−1(4/3)=π/2 and it seems to work for any numbers. Do you know what this identity is?

hero (hero):

If you want integration rules, use these

hero (hero):

From that list, you can apply #8

OpenStudy (anonymous):

I solved it but now the ln part disappears pretty much all together. And I still get the answer of tan-1(3/4) value instead of tan-1(4/3)

OpenStudy (anonymous):

I know how to integrate to get inverse tan functions. But from what you did, the ln part for my integral disappeared all together. Also, as I was saying, I do get the right answer, and also just different tan-1, which is weird. But then this sort of identity thing solves my initial question... I just don't know what it is. But thanks :)